OpenStudy (anonymous):

I need help with dis question! An ant and its cargo weigh 76 milligrams. The cargo is 18 times heavier than the ant. Use a linear system and Cramer's Rule to find the weight of the ant and its cargo.

OpenStudy (triciaal):

let the weight of the ant = x 76 = 18x + x 76 = 19x x = 76/19 = 4 cargo = 18x = 18 (4) = 72 check 4 + 72 = 76

OpenStudy (anonymous):

THanks

OpenStudy (triciaal):

you are more than welcome

OpenStudy (anonymous):

(;